OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{n!}{n^{n}}\] = \[\lim_{ n \rightarrow \infty} \frac{(n+1)!}{(n+1)^{n+1}} \frac{n^{n}}{n!}\] = \[\lim_{ n \rightarrow \infty} \frac{n^{n}}{(n+1)^{n+1}} \frac{(n+1)!}{n!}\] = \[\lim_{ n \rightarrow \infty} \frac{n^{n}}{(n+1)^{n+1}} (n+1)\] = \[\lim_{ n \rightarrow \infty} \frac{n^{n}}{(n+1)^{n}} \] = \[\lim_{ n \rightarrow \infty} \frac{1}{(1+ \frac{1}{n})^{n}} \] = stuck

OpenStudy (turingtest):

exponentiate?

OpenStudy (zarkon):

look at the definition of e

OpenStudy (zarkon):

at least one of the definitions of e

OpenStudy (anonymous):

interesting

OpenStudy (turingtest):

@Zarkon ok, I hate doing it that way, but I should get used to different methods I have seen you, among others, use that way before though, and I do see it

OpenStudy (anonymous):

1/e

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

(1/e) < 1 absolute convergence

OpenStudy (zarkon):

\[\frac{n!}{n^{n}}=\frac{n}{n}\frac{n-1}{n}\cdots\frac{3}{n}\frac{2}{n}\frac{1}{n}\] \[\le\frac{2}{n}\frac{1}{n}=\frac{2}{n^2}\] since \[\frac{n}{n}\frac{n-1}{n}\cdots\frac{3}{n}<1\]

OpenStudy (zarkon):

I like the comparison test for this one

OpenStudy (nenadmatematika):

MathSofiya you're solution is correct...but are you sure that the right way to say is absolute convergence...''absolute'' is mainly used in solving alternating series...just to be sure...maybe in your country you use it anyways...:D

OpenStudy (anonymous):

lolz...according to stewart's calculus definity of the "ratio test"...

OpenStudy (anonymous):

hope it's credible...

OpenStudy (nenadmatematika):

OK,just checking because here in Europe we say convergent or divergent for series which are not alternating...but if converges absolutely then converges..so nevermind :D

OpenStudy (anonymous):

I see. Just out of curiosity...Where in Europe?

OpenStudy (nenadmatematika):

Serbia

OpenStudy (anonymous):

Very Cool!

OpenStudy (anonymous):

More power to Open Study!!! Math students from all over the globe...I just think that's soo cool!!

OpenStudy (turingtest):

OpenStudy is made up of people from all over the world. I personally know someone online from each continent besides Antarctica :)

OpenStudy (anonymous):

Amazing!